### v2

@@ 5.2 Reduction Rules
Definition 6.
Given an instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$, an allocation rule that allocates bundles $A_{1},\ldots,A_{k}$ to agents $1,\ldots,k$ respectively is $\alpha$-valid, if and only if the following holds:
@@
The following lemma with small value of $k$ (namely $k\leq 3$) has been proven and used in [28, 3, 1].
Lemma 8.
Given an ordered fair division instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$, a fixed agent $i\in N$, and integer $k<|M|/n$, let $K=\{k(n-1)+1,k(n-1)+2,\ldots,nk,nk+1\}$. Then, $\mu^{n-1}_{i}(M\setminus K)\geq\mu^{n}_{i}(M).$
Proof.
Let $P=(P_{1},P_{2},\ldots,P_{n})$ be an MMS partition of agent $i$ for the original instance. By the pigeonhole principle, there exists $j\in[n]$ such that $|P_{j}\cap[kn+1]|>k$. Without loss of generality, let us assume $|P_{n}\cap[kn+1]|>k$ and $g_{1},g_{2},\ldots g_{k+1}$ are $k+1$ distinct goods in $P_{n}\cap[kn+1]$ such that $g_{1}\leq g_{2}\leq\ldots\leq g_{k+1}$. For $j\in[k+1]$, let $kn-k+j\in P_{a_{j}}$. Now for all $j\in[k+1]$, swap the goods $kn-k+j$ and $g_{j}$; i.e., iteratively do the following:
Let $P^{\prime}=(P^{\prime}_{1},P^{\prime}_{2},\ldots,P^{\prime}_{n})$ be the final partition. Since $g_{1},\ldots,g_{k+1}$ are $k+1$ goods in $[kn+1]$ in increasing order of index, $g_{j}\leq kn-k+j$. Thus, $v_{i}(g_{j})\geq v_{i}(kn-k+j)$ and after each of these swaps, the value of $P_{a_{j}}$ cannot decrease. Therefore, for all $j\in[n-1]$, $v_{i}(P^{\prime}_{j})\geq v_{i}(P_{j})\geq\mu^{n}_{i}(M)$ and $(P^{\prime}_{1},\ldots,P^{\prime}_{n-1})$ is a partition of a subset of $M\setminus K$ into $n-1$ bundles of value at least $\mu^{n}_{i}(M)$ for $i$. The lemma follows. ∎
Rule $R^{\alpha}_{k}(\mathcal{I})$ Preconditions: $\mathcal{I}$ is ordered. For $K=\{k(n-1)+1,k(n-1)+2,\ldots,nk,nk+1\}$, there exists an agent $i$ such that $v_{i}(K)\geq\alpha\cdot\mu_{i}(\mathcal{I})$. Process: Allocate $K$ to agent $i$. Set $N^{\prime}\leftarrow N\setminus\{i\}$ and $M^{\prime}\leftarrow M\setminus K$. Guarantees: $\mathcal{I}^{\prime}=(N^{\prime},M^{\prime},\{v_{i}\}_{i\in N^{\prime}})$ is ordered. For all $i\in N^{\prime}$, $\mu_{i}(\mathcal{I}^{\prime})\geq\mu_{i}(\mathcal{I})$.
Corollary 4 (of Lemma 8).
Given an ordered instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$, for all $k<|M|/n$ and $\alpha\geq 0$, $R^{\alpha}_{k}(\mathcal{I})$ is an $\alpha$-valid reduction.
Definition 7.
We call an ordered instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$, $R^{\alpha}_{k}$-irreducible, if $R^{\alpha}_{k}(\mathcal{I})$ is not applicable. Moreover, we call an ordered instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$, $R^{\alpha}$-irreducible, if for all $k<|M|/|N|$, $R^{\alpha}_{k}(\mathcal{I})$ is not applicable.
Observation 1.
If $\,\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ is ordered and $R^{\alpha}_{k}$-irreducible for some $k<|M|/n$, then $v_{i}(nk+1)<\alpha\cdot\mu_{i}/(k+1)$ for all agents $i\in N$.
Proof.
@@
Input: An ordered instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ and approximation factor $\alpha$.
@@
Input: An ordered instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ and approximation factor $\alpha$.
@@
Let $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ be a fair division instance with additive valuations, and fix an agent $i\in N$. Let $g_{1},g_{2}\in M$ be such that $v_{i}(g_{1})+v_{i}(g_{2})\leq\mu^{n}_{i}(M)$. Then $\mu^{n}_{i}(M)\leq\mu^{n-1}_{i}(M\setminus\{g_{1},g_{2}\}).$
Proof.
To prove the lemma, it suffices to find a partition of (a subset of) $M\setminus\{g_{1},g_{2}\}$ into $n-1$ bundles, such that the value of agent $i$ for each bundle is at least $\mu^{n}_{i}(M)$. We call such a partition a certificate. Let $P=(P_{1},\ldots,P_{n})$ be an MMS partition of agent $i$ in the original instance $\mathcal{I}$. Without loss of generality, assume $\{g_{1},g_{2}\}\subseteq P_{1}\cup P_{2}$. If $\{g_{1},g_{2}\}\subseteq P_{1}$, then $(P_{2},\ldots,P_{n-1},P_{n})$ is a certificate. Otherwise, let $g_{i}\in P_{i}$ for $i\in[2]$. We have
Lemma 10.
@@
Given a fair division instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ and a fixed agent $i\in N$ with additive valuation, let $g_{1},g_{2},g_{3}\in M$ be such that $g_{2}\neq g_{3}$ and $v_{i}(g_{2})+v_{i}(g_{3})\leq\mu^{n}_{i}(M)$. Then $\mu^{n-2}_{i}(M\setminus\{g_{1},g_{2},g_{3}\})\geq\mu^{n}_{i}(M).$
Proof.
To prove the lemma, it suffices to partition (a subset of) $M\setminus\{g_{1},g_{2},g_{3}\}$ into $n-2$ bundles, each of value at least $\mu^{n}_{i}(M)$ to $i$. We call such a partition a certificate. Let $P=(P_{1},\ldots,P_{n})$ be an MMS partition of agent $i$ in the original instance $\mathcal{I}$. Without loss of generality, assume $\{g_{1},g_{2},g_{3}\}\subseteq P_{1}\cup\ldots\cup P_{j}$ for $j\leq 3$. If $\{g_{1},g_{2},g_{3}\}\subseteq P_{1}\cup P_{2}$, then $(P_{3},\ldots,P_{n-1},P_{n})$ is a certificate. Otherwise, let $g_{i}\in P_{i}$ for $i\in[3]$. We have
@@
Rule $S^{\alpha}(\mathcal{I})$ Preconditions: $\mathcal{I}$ is ordered and $R^{\alpha}_{1}$-irreducible. There exists an agent $i$ such that $v_{i}(\{1,n+1\})\geq\alpha\cdot\mu_{i}(\mathcal{I})$. Process: Allocate $\{1,n+1\}$ to $i$. Set $N^{\prime}\leftarrow N\setminus\{i\}$ and $M^{\prime}\leftarrow M\setminus\{1,n+1\}$. If there exists agent $j\neq i$ such that $v_{i}(\{1,n+1\})\geq\alpha\cdot\mu_{j}(\mathcal{I})$: - Duplicate good $1$ and allocate $\{1,n\}$ to agent $j$. - Set $N^{\prime}\leftarrow N^{\prime}\setminus\{j\}$ and $M^{\prime}\leftarrow M^{\prime}\setminus\{n\}$. Guarantees: $\mathcal{I}^{\prime}=(N^{\prime},M^{\prime},\{v_{i}\}_{i\in N^{\prime}})$ is ordered. For all $i\in N^{\prime}$, $\mu_{i}(\mathcal{I}^{\prime})\geq\mu_{i}(\mathcal{I})$.
Lemma 11.
@@
Given an ordered instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ and $\alpha\leq 1$, $S^{\alpha}(\mathcal{I})$ is an $\alpha$-valid reduction.
Proof.
For all agents $i$ who is assigned a bundle $B$ in $S^{\alpha}$, we have $v_{i}(B)\geq\alpha\cdot\mu_{i}(\mathcal{I})$. First consider the case that $S^{\alpha}$ removes two agents $i\neq j$. Since the reduction rule $R^{\alpha}_{1}$ is not applicable, for all the remaining agents $a$, we have $v_{a}(\{n,n+1\})<\alpha\cdot\mu^{n}_{a}(M)\leq\mu^{n}_{a}(M)$. Hence, by lemma 10, $\mu^{n-2}_{a}(M\setminus\{1,n,n+1\})\geq\mu^{n}_{a}(M)$.
Now consider the case that $S^{\alpha}$ removes only one agent $i$. Let $a\neq i$ be a remaining agent after allocating $\{1,n+1\}$ to $i$. We have $v_{a}(\{1,n+1\})<\alpha\cdot\mu^{n}_{i}(M)\leq\mu^{n}_{i}(M)$, otherwise, $S^{\alpha}$ would allocate $\{1,n\}$ to $a$ (or some agent other than $i$). By Lemma 9, $\mu^{n-1}_{a}(M\setminus\{1,n+1\})\geq\mu^{n}_{a}(M)$. ∎
@@
Rule $T^{\alpha}(\mathcal{I})$ Preconditions: $\mathcal{I}$ is ordered and $R^{\alpha}_{k}$-irreducible for $k\leq 1$. There exists three different agents $i,j,\ell$ such that $v_{i}(\{1,n+1\})\geq\alpha\cdot\mu_{i}(\mathcal{I})$, $v_{j}(\{2,n+2\})\geq\alpha\cdot\mu_{j}(\mathcal{I})$, and $v_{\ell}(\{1,n+3\})\geq\alpha\cdot\mu_{\ell}(\mathcal{I})$. Process: Allocate $\{1,n+1\}$ to $i$. Allocate $\{2,n+2\}$ to $j$. Duplicate good $1$ and allocate $\{1,n+3\}$ to $\ell$. Set $N^{\prime}\leftarrow N\setminus\{i,j,\ell\}$ and $M^{\prime}\leftarrow M\setminus\{1,2,n+1,n+2,n+3\}$. Guarantees: $\mathcal{I}^{\prime}=(N^{\prime},M^{\prime},\{v_{i}\}_{i\in N^{\prime}})$ is ordered. If $\alpha\leq 4/5$, for all $i\in N^{\prime}$, $\mu_{i}(\mathcal{I}^{\prime})\geq\mu_{i}(\mathcal{I})$.
Lemma 12.
@@
To prove the lemma, it suffices to partition (a subset of) $M\setminus S$ into $n-3$ bundles, each of value at least $\mu^{n}_{i}(M)$ to $i$. We call such a partition a certificate. Let $P=(P_{1},\ldots,P_{n})$ be an MMS partition of agent $i$ in the original instance $\mathcal{I}$. Without loss of generality, assume $S\subseteq P_{1}\cup\ldots\cup P_{k}$ for $k\leq 5$. If $S\subseteq P_{1}\cup P_{2}\cup P_{3}$, then $(P_{4},\ldots,P_{n-1},P_{n})$ is a certificate. Otherwise, we consider two cases:
@@
Lemma 13.
Given an ordered instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ and $\alpha\leq 4/5$, $T^{\alpha}(\mathcal{I})$ is an $\alpha$-valid reduction.
@@
For all agents $i$ who is assigned a bundle $B$ in $T^{\alpha}$, we have $v_{i}(B)\geq\alpha\cdot\mu_{i}(\mathcal{I})$. Note that by the preconditions of the reduction $T^{\alpha}(\mathcal{I})$, $\mathcal{I}$ is $R^{\alpha}_{k}$-irreducible for $k\leq 1$. Thus, by Observation 1, for all the remaining agents $i$, we have $\alpha\cdot\mu^{n}_{i}(M)\geq v_{i}(1)\geq v_{i}(2)$ and $\alpha/2\cdot\mu^{n}_{i}(M)\geq v_{i}(n+1)\geq v_{i}(n+2)\geq v_{i}(n+3)$. Since $\alpha\leq 4/5$, by Lemma 12, $v^{n-3}_{i}(M\setminus\{1,2,n+1,n+2,n+3\})\geq v^{n}_{i}(M)$. ∎
Remark 1.
In subsequent sections, similar to Definition 7, an ordered instance is $S^{\alpha}$-irreducible and $T^{\alpha}$-irreducible, if $S^{\alpha}$ and $T^{\alpha}$ are not applicable respectively.
@@ 5.3 Combining Round Robin and Bag-Filling
Note that the input $\mathcal{I}$ of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$ is $R^{\alpha}$-irreducible. For all the agents $i$, we define $\mu_{i}=\mu^{n}_{i}(M)$ to be the MMS value of agent $i$ in the beginning of the Algorithm 4. We do a round robin fashion process as following: starting from $n$ empty bags $B_{1},\ldots,B_{n}$, for $j$ in the cyclic sequence $1,2,\ldots,n$, we add the most valuable available good (not yet allocated to any bag) to $B_{j}$. If for some remaining agent $i$, $v_{i}(B_{j})\geq\alpha\cdot\mu_{i}$, then we allocate $B_{j}$ to $i$ and remove $i$ from the set of agents and $j$ from the cyclic sequence. We call this phase of the algorithm, the round robin phase. Note that for all the remaining bags $B_{a}$ and $B_{b}$, if $a<b$, then $v_{i}(B_{a})\geq v_{i}(B_{b})$ for all agents $i$ and hence, first index $1$ will be removed, then index $2$ and so on (see Lemma 15). Let $\{B_{a},\ldots,B_{n}\}$ be the set of the remaining bags at the end of the round robin phase. In the next phase of the algorithm, called duplication phase, we duplicate goods $11,2,\ldots,n-a+$ 1, add them to $B_{a},\ldots,B_{n}$ respectively and allocate these bags to the remaining agents arbitrarily. Unlike the round robin phase, it is not clear whether the agents who are allocated during the duplication phase value their bundle at least $\alpha$ fraction of their MMS. However, in Sections 5.4 and 5.5, we show that by appropriately reducing the instance in advance, we can ensure both the desired approximation guarantees and the bound on the number of distinct copies.
@@
Input: An $R^{\alpha}$-irreducible instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ and approximation factor $\alpha$
@@
Input: An $R^{\alpha}$-irreducible instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ and approximation factor $\alpha$
@@
Let $g$ be the last good added to $B_{j}$. We have $g\geq 2n+1$ and hence $v_{i}(g)\leq v_{i}(2n+1)\leq\alpha/3$. The last inequality holds since the instance is $R^{\alpha}_{2}$-irreducible (by Observation 1). We have $v_{i}(B_{j}\setminus\{g\})<\alpha$. Otherwise, $B_{j}\setminus\{g\}$ would have been allocated to $i$, or some other agent who values it at least $\alpha$, previously. Therefore we have,
@@
Lemma 15.
Let $\mathcal{B}$ be the set of bags allocated during the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$. Then =$\mathcal{B}=\{B_{1},\ldots,B_{|\mathcal{B}|}\}$.
@@
5.4 $6/7$-MMS with Copies
In this section, we present an algorithm which given an ordered instance, outputs a simple $\alpha$-MMS allocation with at most $\lfloor n/2\rfloor$ distinct copies for any $\alpha\leq 6/7$.
@@ 5.4 $6/7$-MMS with Copies
Input: An ordered instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ and approximation factor $\alpha$
@@
Input: An ordered instance $\mathcal{I}=(N,M,\{v_{i}\}_{i\in N})$ and approximation factor $\alpha$
@@
Lemma 16 follows from Lemma 11 and Corollary 4.
Lemma 16.
@@
For the rest of this (sub)-section, let $\alpha\leq 6/7$, $\mathcal{I}$ be an $R^{\alpha}$-irreducible and $S^{\alpha}$-irreducible instance, and $\mu_{i}(\mathcal{I})=1$. Let $n$ be the number of agents right before the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$, =$\mathcal{B}=\{B_{1},\ldots,B_{a-1}\}$ be the set of bags allocated during this phase, and $n^{\prime}=n-a+1$ be the number of remaining agents after this phase.
Observation 2.
@@
Lemma 17.
Let $i$ be a remaining agent after the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$ (if any). Then |$|\{B\in\mathcal{B}\mid v_{i}(B)>1\}|>n^{\prime}$.
@@
Let +=$\mathcal{B}^{+}=\{B\in\mathcal{B}\mid v_{i}(B)>1\}$. By definition, for all $B\in\mathcal{B}\setminus\mathcal{B}^{+}$, $v_{i}(B)\leq 1$. By Observation 2, for all $B\in\mathcal{B}$ , $|B|\geq 3$. By Lemma 14, for all $B\in\mathcal{B}^{+}$, $v_{i}(B)\leq 4\alpha/3$. Also, let $\bar{\mathcal{B}}$ be the set of $n-|\mathcal{B}|=n^{\prime}$ remaining bags after the round robin phase. For all $B\in\bar{\mathcal{B}}$, $v_{i}(B)<\alpha$. Otherwise, $B$ should have been allocated to $i$ (or some other agent who values it at least $\alpha$). Therefore, we have
@@
Removing $(|\mathcal{B}|-|\mathcal{B}^{+}|)$ from both sides of the inequality, we obtain,

### v3

@@ 4.2 Reduction Rules
Definition 4.8.
Given an instance $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$, an allocation rule that allocates bundles $A_{1},\ldots,A_{k}$ to agents $1,\ldots,k$ respectively is $\alpha$-valid, if and only if the following holds:
@@
The following lemma with small value of $k$ (namely $k\leq 3$) has been proven and used in [garg2021improved, akrami2023simplification, akrami2023breaking34barrierapproximate].
Lemma 4.9.
Given an ordered fair division instance $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$, a fixed agent $i\in N$, and integer $k<|M|/n$, let $K=\{k(n-1)+1,k(n-1)+2,\ldots,nk,nk+1\}$. Then, $\mu^{n-1}_{i}(M\setminus K)\geq\mu^{n}_{i}(M).$
Rule $R^{\alpha}_{k}(\mathcal{I})$ Preconditions: For $K=\{k(n-1)+1,\ldots,nk,nk+1\}$, there exists an agent $i$ such that $v_{i}(K)\geq\alpha\cdot\mu_{i}(\mathcal{I})$. Process: Allocate $K$ to agent $i$. Set $N^{\prime}\leftarrow N\setminus\{i\}$ and $M^{\prime}\leftarrow M\setminus K$. Guarantees: For all $i\in N^{\prime}$, $\mu_{i}(\mathcal{I}^{\prime})\geq\mu_{i}(\mathcal{I})$.
Corollary 4.10 (of Lemma 4.9).
Given an ordered instance $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$, for all $k<|M|/n$ and $\alpha\geq 0$, $R^{\alpha}_{k}(\mathcal{I})$ is an $\alpha$-valid reduction.
Definition 4.11.
We call an ordered instance $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$, $R^{\alpha}_{k}$-irreducible, if $R^{\alpha}_{k}(\mathcal{I})$ is not applicable. Moreover, we call an ordered instance $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$, $R^{\alpha}$-irreducible, if for all $k<|M|/|N|$, $R^{\alpha}_{k}(\mathcal{I})$ is not applicable.
Observation 4.12.
If $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$ is ordered and $R^{\alpha}_{k}$-irreducible for some $k<|M|/n$, then $v_{i}(nk+1)<\alpha\cdot\mu_{i}/(k+1)$ for all agents $i\in N$.
@@
Input: An ordered instance $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$ and approximation factor $\alpha$.
@@
Input: An ordered instance $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$ and approximation factor $\alpha$.
@@
Let $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$ be a fair division instance with additive valuations, and fix an agent $i\in N$. Let $g_{1},g_{2}\in M$ be such that $v_{i}(g_{1})+v_{i}(g_{2})\leq\mu^{n}_{i}(M)$. Then $\mu^{n}_{i}(M)\leq\mu^{n-1}_{i}(M\setminus\{g_{1},g_{2}\}).$
Lemma 4.14.
@@
Given a fair division instance $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$ and a fixed agent $i\in N$ with additive valuation, let $g_{1},g_{2},g_{3}\in M$ be such that $g_{2}\neq g_{3}$ and $v_{i}(g_{2})+v_{i}(g_{3})\leq\mu^{n}_{i}(M)$. Then $\mu^{n-2}_{i}(M\setminus\{g_{1},g_{2},g_{3}\})\geq\mu^{n}_{i}(M).$
@@
Rule $S^{\alpha}(\mathcal{I})$ Preconditions: $\mathcal{I}$ is $R^{\alpha}_{1}$-irreducible. There exists an agent $i$ such that $v_{i}(\{1,n+1\})\geq\alpha\cdot\mu_{i}(\mathcal{I})$. Process: Allocate $\{1,n+1\}$ to $i$. Set $N^{\prime}\leftarrow N\setminus\{i\}$ and $M^{\prime}\leftarrow M\setminus\{1,n+1\}$. If there exists agent $j\neq i$ such that $v_{j}(\{1,n+1\})\geq\alpha\cdot\mu_{j}(\mathcal{I})$: - Duplicate good $1$ and allocate $\{1,n\}$ to agent $j$. - Set $N^{\prime}\leftarrow N^{\prime}\setminus\{j\}$ and $M^{\prime}\leftarrow M^{\prime}\setminus\{n\}$. Guarantees: For all $i\in N^{\prime}$, $\mu_{i}(\mathcal{I}^{\prime})\geq\mu_{i}(\mathcal{I})$.
Lemma 4.15.
@@
Given an ordered instance $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$ and $\alpha\leq 1$, $S^{\alpha}(\mathcal{I})$ is an $\alpha$-valid reduction.
@@
Rule $T^{\alpha}(\mathcal{I})$ Preconditions: $\mathcal{I}$ is $R^{\alpha}_{k}$-irreducible for $k\leq 1$. There exists three different agents $i,j,\ell$ such that $v_{i}(\{1,n+1\})\geq\alpha\cdot\mu_{i}(\mathcal{I})$, $v_{j}(\{2,n+2\})\geq\alpha\cdot\mu_{j}(\mathcal{I})$, and $v_{\ell}(\{1,n+3\})\geq\alpha\cdot\mu_{\ell}(\mathcal{I})$. Process: Allocate $\{1,n+1\}$ to $i$ and $\{2,n+2\}$ to $j$. Duplicate good $1$ and allocate $\{1,n+3\}$ to $\ell$. Set $N^{\prime}\leftarrow N\setminus\{i,j,\ell\}$ and $M^{\prime}\leftarrow M\setminus\{1,2,n+1,n+2,n+3\}$. Guarantees: If $\alpha\leq 4/5$, for all $i\in N^{\prime}$, $\mu_{i}(\mathcal{I}^{\prime})\geq\mu_{i}(\mathcal{I})$.
Lemma 4.16.
@@
To prove the lemma, it suffices to partition (a subset of) $M\setminus S$ into n3n-3 bundles, each of value at least $\mu^{n}_{i}(M)$ to $i$. We call such a partition a certificate. Let $P=(P_{1},\ldots,P_{n})$ be an MMS partition of agent $i$ in the original instance $\mathcal{I}$. Without loss of generality, assume $S\subseteq P_{1}\cup\ldots\cup P_{k}$ for $k\leq 5$. If $S\subseteq P_{1}\cup P_{2}\cup P_{3}$, then $(P_{4},\ldots,P_{n-1},P_{n})$ is a certificate. Otherwise, we consider two cases:
@@
Lemma 4.17.
Given an ordered instance $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$ and $\alpha\leq 4/5$, $T^{\alpha}(\mathcal{I})$ is an $\alpha$-valid reduction.
@@
For all agents $i$ who is assigned a bundle $B$ in $T^{\alpha}$, we have $v_{i}(B)\geq\alpha\cdot\mu_{i}(\mathcal{I})$. Note that by the preconditions of the reduction $T^{\alpha}(\mathcal{I})$, $\mathcal{I}$ is $R^{\alpha}_{k}$-irreducible for $k\leq 1$. Thus, by Observation 4.12, for all the remaining agents $i$, we have $\alpha\cdot\mu^{n}_{i}(M)\geq v_{i}(1)\geq v_{i}(2)$ and $\alpha/2\cdot\mu^{n}_{i}(M)\geq v_{i}(n+1)\geq v_{i}(n+2)\geq v_{i}(n+3)$. Since $\alpha\leq 4/5$, by Lemma 4.16, $v^{n-3}_{i}(M\setminus\{1,2,n+1,n+2,n+3\})\geq v^{n}_{i}(M)$. ∎
Remark.
In subsequent sections, similar to Definition 4.11, an ordered instance is $S^{\alpha}$-irreducible and $T^{\alpha}$-irreducible, if $S^{\alpha}$ and $T^{\alpha}$ are not applicable respectively.
@@ 4.3 Combining Round Robin and Bag-Filling
Note that the input $\mathcal{I}$ of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$ is $R^{\alpha}$-irreducible. For all the agents $i$, we define $\mu_{i}=\mu^{n}_{i}(M)$ to be the MMS value of agent $i$ in the beginning of the Algorithm 4. We do a round robin fashion process as following: starting from $n$ empty bags $B_{1},\ldots,B_{n}$, for $j$ in the cyclic sequence $1,2,\ldots,n$, we add the most valuable available good (not yet allocated to any bag) to $B_{j}$. If for some remaining agent $i$, $v_{i}(B_{j})\geq\alpha\cdot\mu_{i}$, then we allocate $B_{j}$ to $i$ and remove $i$ from the set of agents and $j$ from the cyclic sequence. We call this phase of the algorithm, the round robin phase. Note that for all the remaining bags $B_{a}$ and $B_{b}$, if $a<b$, then $v_{i}(B_{a})\geq v_{i}(B_{b})$ for all agents $i$ and hence, first index $1$ will be removed, then index $2$ and so on (see Lemma 4.19). Let $\{B_{a},\ldots,B_{n}\}$ be the set of the remaining bags at the end of the round robin phase. In the next phase of the algorithm, called duplication phase, we duplicate goods $1,2,\ldots,n-a+1$, add them to $B_{a},\ldots,B_{n}$ respectively and allocate these bags to the remaining agents arbitrarily. Unlike the round robin phase, it is not clear whether the agents who are allocated during the duplication phase value their bundle at least $\alpha$ fraction of their MMS. However, in Sections 4.4 and 4.5, we show that by appropriately reducing the instance in advance, we can ensure both the desired approximation guarantees and the bound on the number of distinct copies.
@@
Input: An $R^{\alpha}$-irreducible instance $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$ and approximation factor $\alpha$
@@
Input: An $R^{\alpha}$-irreducible instance $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$ and approximation factor $\alpha$
@@
Let $g$ be the last good added to $B_{j}$. We have $g\geq 2n+1$ and hence $v_{i}(g)\leq v_{i}(2n+1)\leq\alpha/3$. The last inequality holds since the instance is $R^{\alpha}_{2}$-irreducible (by Observation 4.12). We have $v_{i}(B_{j}\setminus\{g\})<\alpha$. Otherwise, $B_{j}\setminus\{g\}$ would have been allocated to $i$, or some other agent who values it at least $\alpha$, previously. Therefore we have,
@@
Lemma 4.19.
Let $\mathcal{B}$ be the set of bags allocated during the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$. Then $\mathcal{B}=\{B_{1},\ldots,B_{|\mathcal{B}|}\}$.
@@
[…]
In this section, we present an algorithm which given an ordered instance, outputs a simple $\alpha$-MMS allocation with at most $\lfloor n/2\rfloor$ distinct copies for any $\alpha\leq 6/7$. Then, using Proposition 4.1, we prove the following theorem.
See 4.2
@@
Input: An ordered instance $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$ and approximation factor $\alpha$
@@
Input: An ordered instance $\mathcal{I}=\langle N,M,\{v_{i}\}_{i\in N}\rangle$ and approximation factor $\alpha$
@@
Lemma 4.20 follows from Lemma 4.15 and Corollary 4.10.
Lemma 4.20.
@@
For the rest of this (sub)-section, let $\alpha\leq 6/7$, $\mathcal{I}$ be an $R^{\alpha}$-irreducible and $S^{\alpha}$-irreducible instance, and $\mu_{i}(\mathcal{I})=1$. Let $n$ be the number of agents right before the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$, $\mathcal{B}=\{B_{1},\ldots,B_{a-1}\}$ be the set of bags allocated during this phase, and $n^{\prime}=n-a+1$ be the number of remaining agents after this phase.
Observation 4.21.
@@
Lemma 4.22.
Let $i$ be a remaining agent after the round robin phase of $\mathsf{BagFill\mbox{-}RoundRobin}(\mathcal{I},\alpha)$ (if any). Then $|\{B\in\mathcal{B}\mid v_{i}(B)>1\}|>n^{\prime}$.
@@
Let $\mathcal{B}^{+}=\{B\in\mathcal{B}\mid v_{i}(B)>1\}$. By definition, for all $B\in\mathcal{B}\setminus\mathcal{B}^{+}$, $v_{i}(B)\leq 1$. By Observation 4.21, for all $B\in\mathcal{B}$ , $|B|\geq 3$. By Lemma 4.18, for all $B\in\mathcal{B}^{+}$, $v_{i}(B)\leq 4\alpha/3$. Also, let $\bar{\mathcal{B}}$ be the set of $n-|\mathcal{B}|=n^{\prime}$ remaining bags after the round robin phase. For all $B\in\bar{\mathcal{B}}$, $v_{i}(B)<\alpha$. Otherwise, $B$ should have been allocated to $i$ (or some other agent who values it at least $\alpha$). Therefore, we have
@@
Removing (|||+|)($\mathcal{B}|$-$\mathcal{B}^{+}|$) from both sides of the inequality, we obtain,